A classic spring/block system is moving back and forth on an air track (no friction). The mass of the block is 0.5 kg and the spring constant is 200 N/m. The mass is pulled back from equilibrium and then quickly pushed into motion such that it has an initial speed at the release point (60 cm from the equilibrium position) of 5.0 m/s. Determine the frequency of this oscillator.
Answer #1

Mass of block,m=0.5 kg

Spring constant,k=200 N/m

Frequency of oscillator is given by

f=(1/2\pi) (k/m)1/2

=(1/2\pi) (200/0.5)1/2=3.18 Hz

Frequency of this oscillator.=3.18 Hz